# Random close packing of binary hard spheres predicts the stability of atomic nuclei

### v1

@@ Abstract
In spite of the empirical success of the Weizsäcker mass formula Weizsäcker (1935); Mun et al. (2024), a predictive, physics-based, analytical estimate of the stability chart of nuclei has remained elusive. We show that a recent analytical progress in the mathematical description of random close packing of spheres with different sizes Anzivino et al. (2023) provides an excellent description of the $Z$ versus $N$ slope in the nuclides chart. In particular, the theory is able to predict the most quoted ratio $Z/N\approx 0.75$ for stable nuclides Preston (1962), in agreement with experimental observations without any fitting parameter. This solves a 90-year old mystery in nuclear physics: until now, the stability of atomic nuclei was empirically described by the Bethe-Weizsäcker mass formula, which, only thanks to 4-5 adjustable fitting parameters Segrè (197), can capture the tendency of large nuclei to be rich in neutrons. The latter phenomenon was traditionally (qualitatively) explained in terms of a subtle balance between electrostatics and asymmetry effects due to the Pauli exclusion principle. We show that the larger stability of neutron-rich nuclei is the result of a maximized nuclear density, which is controlled by geometric packing considerations.
@@
where $A=Z+N$. The first term on the right hand side is the so-called volume term due to nucleon-nucleon attractive interaction energy, the second term is the surface term, the third one is the Coulomb repulsion term due to the protons, the fourth term is the asymmetry term, and the last one is the pairing term. The respective parameters ($a_{V}$, $a_{S}$, $a_{C}$, $a_{A}$, $\delta$) are empirical coefficients, the values of which are determined by fitting experimental data from mass spectroscopy Bohr and Mottelson (1998). Of these 5 fitting parameters, perhaps only the Coulomb term can be precisely determined from theory. More sophisticated versions of Eq. (1) exist Moller et al. (1995), which contain even more parameters. With 4-5 adjustable parameters it is no mystery that Eq. (1) can be used to provide an excellent fitting to both the binding energy curve of all nuclides and to the Segrè chart of stability of nuclides (where for each nuclide the proton number $Z$ is plotted as a function of the neutron number $N$) Segrè (197).
@@
Given the lack of predictive power of the Weizsäcker mass formula Eq. (1), it is currently impossible to quantitatively understand or predict the Segrè chart of nuclear stability. As is well known, Eq. (1) shows the tendency to predict that the most stable nuclei are those with $Z=N$. Only thanks to the empirical fitting, the adjustment of the parameters to the experimental data can force Eq. (1) to capture the actual observed trend, which has a lower slope than 1, i.e. $Z\approx 0.75N$ Marmier and Sheldon (1969). The standard qualitative understanding of the Segrè nuclear stability chart based on Eq. (1) is that the Coulomb repulsion, for large nuclei, grows as $\sim Z^{2}$, whereas the attractive energy due to the strong force increases only as $\sim Z$ (due to its short-range character). This fact justifies the tendency of larger nuclei to have more neutrons than protons, $N>Z$, in order to compensate the unfavourable growth of the Coulomb energy Segrè (197). However, this energetic gain is, in its own turn, compensated or reduced by the asymmetry between neutrons and protons, which corresponds to the energetically unfavourable asymmetry term, i.e. the fourth term in Eq. (1). In a nutshell, this term arises due to the Pauli exclusion principle, because having more neutrons than protons ($N>Z)$ implies that the neutrons have to be stacked up into higher energy levels. More specifically, there are two contributions from the asymmetry (neutron excess). The first is an increase in the total potential energy, which can be qualitatively estimated by considering that the nuclear wavefunction is anti-symmetric with respect to exchanging two identical particles in different space states. The energy cost of interchanging protons with neutrons can be evaluated by making use of the Majorana space-exchange forces, leading to a contribution to the binding of the form $\sim(N-Z)^{2}/A$ Blatt and Weisskopf (1979); Preston (1962). The corresponding increase in kinetic energy can be roughly evaluated using a Fermi gas model and also contributes a term of the form $\sim(N-Z)^{2}/A$ Blatt and Weisskopf (1979); Preston (1962). Combining the two contributions with a common prefactor leads to the fourth term on the r.h.s. of Eq. (1). The prefactor $a_{c}$ is necessarily empirical since it is prohibitively difficult to estimate it from theory for a large number of nuclides, and would necessarily be inaccurate given the approximate character of the theory of exchange forces in nuclei Obertelli and Sagawa (2021).
@@
In the following, we apply the theory of RCP for binary mixtures to atomic nuclei and show that it is able to predict the stability of atomic nuclei. We start from the derivation of the RCP for binary mixtures of hard spheres with two different diameters, $\sigma_{1}$ and $\sigma_{2}$, representing, respectively, the diameter of neutrons and protons. The full details of the derivation can be found in previous work Anzivino et al. (2023), here we limit ourselves to a brief recap.
@@
We now consider the following estimates for the proton radius and for the neutron radius. For the proton, the internationally accepted value of its radius is $\sigma_{2}/2\equiv r_{Z}=0.84$ fm NIS . For the neutron, there is no such clear cut value, given the well known difficulties in experimentally determining the size of electrically neutral particles. While the charge radius is not a representative measure since it is based on electrostatics and thus significantly underestimates the actual neutron radius, a more sensible choice is to take the size at which the neutron charge density profile goes to zero, which is at $\sigma_{1}\equiv r_{N}\approx 1.0$ fm according to the latest literature Atac et al. (2021).
@@
By analytically solving Eq. (3) for $\phi$ with $\sigma_{2}/\sigma_{1}=0.84$ we thus obtain the RCP volume fractions plotted in Fig. 1(a) as a function of the ratio of small spheres over the total (in our case, the ratio of protons to the total number of nucleons in the nucleus). We also show in Fig. 1(b) an artistic rendering of the RCP structure with protons and neutrons having the sizes and composition at which the theory predicts the largest packing density (i.e. the maximum of the curve in Fig. 1(a)). For illustrative purposes we chose a simple spherical shape, but it is important to point out that the predicted RCP volume fraction is independent of the nucleus shape. That is, the model predicts the same maximal RCP volume fraction value (for $Z/A=0.43$) also for different shapes, e.g. ellipsoidal or elongated nuclei, and deformed nuclei.
Figure 1: Panel (a) shows the random close packing (RCP) volume fractions computed, for binary hard sphere mixtures where the small spheres have the radius of the proton and the big spheres have the radius of the neutron, based on Eqs. (3)-(5). The theory predicts a maximum RCP density corresponding to the composition $Z/A=0.43$ (dashed vertical line). Panel (b) shows a schematic rendering of an RCP structure corresponding to the largest packing fraction predicted by the theory (indicated by an arrow in panel (a)), i.e. with $Z/A=0.43$ and with size ratio $r_{Z}/r_{N}=\sigma_{2}/\sigma_{1}=0.84$. Red spheres are neutrons and blue spheres are protons. For simplicity, no effects like neutron skin or $\alpha$-particle clustering are taken into account in the picture. The spherical shape was chosen for illustrative purposes, and the predicted RCP volume fraction would be exactly the same also for different nuclear shapes (e.g. ellipsoidal, elongated etc).
@@
From a physical point of view, the above theory explains that the neutron excess experimentally observed in stable nuclei, besides the subtle balance between electrostatics and Pauli-exclusion asymmetry energy, can be understood in terms of the tendency of the nuclei to minimize their energy by finding the configuration with the largest volume fraction occupied by the nucleons, i.e. with the highest density of nucleons. This can be further understood as follows. Due to the short range character of the strong force, the central attractive interaction between nucleons becomes negligible at a nucleon-nucleon separation of about 3 fm Epelbaum et al. (2009, 2020). This implies that only nearest-neighbour and next nearest-neighbour interactions contribute to the binding energy Obertelli and Sagawa (2021). If the nearest-neighbour coordination number is $z$, the total number of pairwise nucleon-nucleon interactions in the nucleus is $zA/2$ (which justifies the first term in Eq. (1) being $\propto A$). If each nucleon has, on average, $z$ nearest neighbours and $z_{nn}$ next nearest neighbours, the total number of interactions contributing to the binding energy is $z_{tot}A/2=(z+z_{nn})A/2$. In the RCP of binary spheres, the total average $z$ is always 6 regardless of the composition i.e. regardless of $Z/A$. However, the mean number of next nearest neighbours is, instead, larger the larger the volume fraction $\phi_{RCP}$. This can be shown more quantitatively by recalling that $z_{tot}$ is given by:
@@
In summary, we have presented a theory of the stability of atomic nuclei in terms of random close packing of binary hard spheres. The theory shows that the binding energy is maximized by the largest nucleons packing, which is estimated using the statistical mechanics theory of random close packing (RCP) of bi-disperse hard spheres Anzivino et al. (2023). Assuming reasonable values for the proton radius and for the neutron radius, the theory predicts that the largest density is achieved when $Z/N=0.75$, the most quoted slope value in the literature Preston (1962). This is the first prediction (with no adjustable parameters) of the observed slope in the experimental table of stable nuclides (the Segrè chart). It comes almost 90 years after the semi-empirical Bethe-Weizsäcker mass formula, which, even in its most sophisticated versions, is unable to predict the correct trend (without using 4-5 adjustable fitting parameters). This result opens up new avenues in low energy nuclear physics. Most importantly, it provides a new rational and predictive guideline in the experimental search for new heavy nuclei and the stability thereof Block et al. (2010); Morjean et al. (2008). It may also have broader implications for nuclear reactions such as fission Wilson et al. (2021) and fusion-fission Eccles et al. (2017), for the equation of state of nuclear matter et al. (2024); Burgio and Vidaña (2020), for the viscosity and hydrodynamics of nuclear matter Grefa et al. (2022); Zaccone (2024), and nucleosynthesis Sun et al. (2024). In future work, the proposed framework may also be extended to deformed warm nuclei to describe dipole oscillations Bortignon et al. (1998).

### v4

# Random close packing of binary hard spheres favors the stability of neutron-rich atomic nuclei

@@ Abstract
In spite of the success of the Bethe-Weizsäcker mass formula in its modern numerical and predictive implementations, the common-knowledge principle that it is electrostatics which, ultimately, favors neutron-rich nuclei still presents unclear aspects. For example, while it is true that the Coulomb interaction promotes the tendency towards neutron-rich nuclei, the opposite effects of Majorana exchange forces and Pauli exclusion are known to counteract this tendency. We show that a recent analytical progress in the mathematical description of random close packing of spheres with different sizes provides a missing contribution to the theoretical description of the $Z$ versus $N$ slope in the nuclides chart. In particular, the theory suggests, on geometric grounds and with a physically-reasoned assumption that the excluded-volume size of neutrons is 20% larger than that of protons, that the most stable nuclei are those with ratio $Z/N\approx 0.75$. This new “geometric” random-packing contribution to the semi-empirical mass formula may be the missing aspect of nuclear structure that tilts the balance towards neutron-rich nuclei in the Segrè stability chart.
@@
where $A=Z+N$. The first term on the right hand side is the so-called volume term due to nucleon-nucleon attractive interaction energy, the second term is the surface term, the third one is the Coulomb repulsion term due to the protons, the fourth term is the asymmetry term, and the last one is the pairing term. The respective parameters ($a_{V}$, $a_{S}$, $a_{C}$, $a_{A}$, $\delta$) are empirical coefficients, the values of which are determined by fitting experimental data from mass spectroscopy Bohr and Mottelson (1998). Of these 5 fitting parameters, perhaps only the Coulomb term can be precisely determined from theory. More sophisticated versions of Eq. (1) exist Moller et al. (1995), as nuclear structure theory has developed into a field of systematic calculations and predictions based on ab initio few- and many-body calculations, reaching nuclei as heavy as 208Pb, as well as based on global energy density functional calculations of the nuclear stability chart. With these auxiliary microscopic calculation methods, Eq. (1) can be used to provide an excellent fitting to both the binding energy curve of all nuclides and to the Segrè chart of stability of nuclides (where for each nuclide the proton number $Z$ is plotted as a function of the neutron number $N$) Segrè (197).
In spite of this, in an attempt to extract general physical principles, it is still interesting to study Eq. (1) from the point of view of how the interplay between different physical phenomena leads Eq. (1) to capture the actual observed stability trend, i.e. $Z\approx 0.75N$ Marmier and Sheldon (1969).
@@
The standard qualitative understanding of the Segrè nuclear stability chart based on Eq. (1) is that the Coulomb repulsion, for large nuclei, grows as $\sim Z^{2}$, whereas the attractive energy due to the strong force increases only as $\sim Z$ (due to its short-range character). This fact justifies the tendency of larger nuclei to have more neutrons than protons, $N>Z$, in order to compensate the unfavourable growth of the Coulomb energy Segrè (197). However, this energetic gain is, in its own turn, compensated or reduced by the asymmetry between neutrons and protons, which corresponds to the energetically unfavourable asymmetry term, i.e. the fourth term in Eq. (1). In a nutshell, this term arises due to the Pauli exclusion principle, because having more neutrons than protons ($N>Z)$ implies that the neutrons have to be stacked up into higher energy levels. More specifically, there are two contributions from the asymmetry (neutron excess). The first is an increase in the total potential energy, which can be qualitatively estimated by considering that the nuclear wavefunction is anti-symmetric with respect to exchanging two identical particles in different space states. The energy cost of interchanging protons with neutrons can be approximately evaluated by making use of the Majorana space-exchange forces, leading to a contribution to the binding of the form $\sim(N-Z)^{2}/A$ Blatt and Weisskopf (1979); Preston (1962). The corresponding increase in kinetic energy can be evaluated, in a rudimentary way, using a Fermi gas model and also contributes a term of the form $\sim(N-Z)^{2}/A$ Blatt and Weisskopf (1979); Preston (1962). Combining the two contributions with a common prefactor leads to the fourth term on the r.h.s. of Eq. (1). The prefactor $a_{c}$ is necessarily empirical since it is prohibitively difficult to estimate it from theory for a large number of nuclides (and it would necessarily be inaccurate given the approximate character of the theory of exchange forces in nuclei) Obertelli and Sagawa (2021).
@@
Here, our focus is not to make accurate quantitative predictions of nuclear stability, which is a task that is best achieved using the current semi-empirical numerical packages Mun et al. (2024). Rather, we want to address an aspect of nuclear stability that has been neglected so far. Common knowledge identifies the main source of isospin breaking in nuclei with the long-range Coulomb interaction, with possibly a subsidiary role played by quark masses. It cannot be denied, however, that the effect of the Coulomb interaction may be compensated or cancelled out by the opposite effect of the exchange forces and by the Pauli principle, as discussed above. Given the presence of empirical parameters, state-of-art numerical calculations cannot provide transparent insights into this question. While keeping the analysis at a fundamental level, we want to show that there is a further structural aspect, neglected in the literature so far, which effectively helps the Coulomb repulsion to win the balance towards favoring neutron-rich nuclei. This effect is purely structural and has to do with the maximal packing density that a mixture of spheres with two different sizes can achieve. This aspect is not in contradiction or in opposition with all the other (mostly quantum-mechanical) aspects of nuclear structure recalled above, and it would represent an additional, new contribution to the semi-empirical mass formula, acting in synergy with the known ones to produce the observed stability trend.
@@
In the following, we apply the theory of RCP for binary mixtures to atomic nuclei and show that it is able to predict the correct stability trend of atomic nuclei. We start from the derivation of the RCP for binary mixtures of hard spheres with two different diameters, $\sigma_{1}$ and $\sigma_{2}$, representing, respectively, the diameter of neutrons and protons. The full details of the derivation can be found in previous work Anzivino et al. (2023), here we limit ourselves to a brief recap.
@@
We now consider the following estimates for the proton radius and for the neutron radius. For the proton, the internationally accepted value of its radius is $\sigma_{2}/2\equiv r_{Z}=0.84$ fm NIS . For the neutron, there is no such clear cut value, given the well known difficulties in experimentally determining the size of electrically neutral particles. It would not make sense to use the root mean square charge radius here since it is related to the neutron-electron scattering length, hence to how neutrons interact with electrons, which is certainly irrelevant and misleading for our purpose of modelling excluded-volume interactions between nucleons in a densely packed nucleus. A more meaningful choice for our purpose of modelling excluded-volume between nucleons is to take the size at which the neutron charge density profile goes to zero, which is at $\sigma_{1}\equiv r_{N}\approx 1.0$ fm, according to the latest literature Atac et al. (2021). This is of course, a hypothesis of the current calculation, which should be granted the benefit of doubt, but it does seem to be the most reasonable one given the alternatives at hand and is not disproved by the available empirical evidence (here one should also be aware of the current unsettled debate on the sizes of the nucleons, especially given the ambiguities which plague the precise determination of charge density distributions of nucleons Jaffe (2021); Epelbaum et al. (2022)).
@@
By analytically solving Eq. (3) for $\phi$ with $\sigma_{2}/\sigma_{1}=0.84$, we thus obtain the RCP volume fractions plotted in Fig. 1(a) as a function of the ratio of small spheres over the total (in our case, the ratio of protons to the total number of nucleons in the nucleus). We also show, in Fig. 1(b), an artistic rendering of the RCP structure with protons and neutrons having the sizes and composition at which the theory predicts the largest packing density (i.e. the maximum of the curve in Fig. 1(a)). For illustrative purposes we chose a simple spherical shape, but it is important to point out that the predicted RCP volume fraction is independent of the nucleus shape. That is, the model predicts the same maximal RCP volume fraction value (for $Z/A=0.43$) also for different shapes, e.g. ellipsoidal or elongated nuclei, and deformed nuclei.
Figure 1: Panel (a) shows the random close packing (RCP) volume fractions computed, for binary hard sphere mixtures where the small spheres have the radius of the proton and the big spheres have the radius of the neutron, based on Eqs. (3)-(5). The theory predicts a maximum RCP density corresponding to the composition $Z/A=0.43$ (dashed vertical line). Panel (b) shows a schematic rendering of an RCP structure corresponding to the largest packing fraction predicted by the theory (indicated by a dashed vertical line in panel (a)), i.e. with $Z/A=0.43$ and with size ratio $r_{Z}/r_{N}=\sigma_{2}/\sigma_{1}=0.84$. Red spheres are neutrons and blue spheres are protons. For simplicity, no effects like neutron skin or $\alpha$-particle clustering are taken into account in the picture. The spherical shape was chosen for illustrative purposes, and the predicted RCP volume fraction would be exactly the same also for different nuclear shapes (e.g. ellipsoidal, elongated etc).
@@
From a physical point of view, the above theory suggests that the neutron-excess observed in stable large nuclei, besides the subtle balance between electrostatics and Pauli-exclusion asymmetry energy, can be understood in terms of the tendency of the nuclei to minimize their energy by finding the configuration with the largest volume fraction occupied by the nucleons, i.e. with the highest density of nucleons. This can be further understood as follows. Due to the short-range character of the strong force, the central attractive interaction between nucleons becomes negligible at a nucleon-nucleon separation of about 3 fm Epelbaum et al. (2009, 2020). This implies that mostly nearest-neighbour and next nearest-neighbour interactions contribute to the binding energy, which is, after all, textbook knowledge Obertelli and Sagawa (2021). If the nearest-neighbour coordination number is $z$, the total number of pairwise nucleon-nucleon interactions in the nucleus is $zA/2$ (which justifies the first term in Eq. (1) being $\propto A$). If each nucleon has, on average, $z$ nearest neighbours and $z_{nn}$ next nearest neighbours, the total number of interactions contributing to the binding energy is $z_{tot}A/2=(z+z_{nn})A/2$. In the RCP of binary spheres, the total average $z$ is always 6 regardless of the composition i.e. regardless of $Z/A$. However, the mean number of next nearest neighbours is, instead, larger the larger the volume fraction $\phi_{RCP}$. This can be shown more quantitatively by recalling that $z_{tot}$ is given by:
@@
In summary, we have presented a contribution to the theory of the stability of atomic nuclei in terms of random close packing of binary hard spheres. The theory shows that the binding energy is maximized by the largest nucleons packing density, which is estimated using the statistical mechanics theory of random close packing (RCP) of bi-disperse hard spheres Anzivino et al. (2023). Assuming reasonable values for the proton radius and for the neutron radius, the theory predicts that the largest density is achieved when $Z/N=0.75$, the most quoted slope value in the literature Preston (1962).
This result opens up new possibilities in nuclear structure and in low energy nuclear physics. Most importantly, it provides a new rational and predictive guideline in the experimental search for new heavy nuclei and the stability thereof Block et al. (2010); Morjean et al. (2008). It may also have broader implications for nuclear reactions such as fission Wilson et al. (2021) and fusion-fission Eccles et al. (2017), for the equation of state of nuclear matter et al. (2024); Burgio and Vidaña (2020), for the viscosity and hydrodynamics of nuclear matter Grefa et al. (2022); Zaccone (2024), and nucleosynthesis Sun et al. (2024). In future work, the proposed ideas may also be extended to deformed warm nuclei to describe dipole oscillations Bortignon et al. (1998).